OpenStudy (kaylak12345):

i need help y=-0.04x^2+1.3x+56 @SolomonZelman

OpenStudy (anonymous):

hold on one moment and ill see if i can solve pk :)

OpenStudy (kaylak12345):

step by step plz

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Do you want the graph or the answer to y and x? What are you really asking for? Just to solve?

OpenStudy (anonymous):

Simplifying y = -0.04x2 + 1.3x + 56 Reorder the terms: y = 56 + 1.3x + -0.04x2 Solving y = 56 + 1.3x + -0.04x2 Solving for variable 'y'. Move all terms containing y to the left, all other terms to the right. Simplifying y = 56 + 1.3x + -0.04x2 Is this what you're looking for?

OpenStudy (kaylak12345):

no a quadratic answer the answer is 57.04 but im not sure how I got it

OpenStudy (anonymous):

Oh ok :) You got it by simplifying the equation to the point where you could solve right?

OpenStudy (kaylak12345):

i guess all i know is i did the quadratic formula

OpenStudy (anonymous):

Ok Well then I guess that it :) However you figured it out is best for you bust yes usually you just simplify and solve.

OpenStudy (kaylak12345):

thats what i do but im told to use stupid formulass